OpenStudy (savannah_noelle):

Can someone please check my factoring homework? I'm not sure if I did everything correctly. I'll give a medal :) 1. Factor: X^2+9X+20 Answer: (X+5)(X+4) 2. Factor: X^2-14X+40 Answer: (X-4)(X-10) 3. Factor: X^2+6X+3 Answer: (X+2)(X+3) 4.Factor: X^2-5X-50 Answer: (X+5)*X-10) 5. Factor: 3M^2+MT-10T^2 Answer: (M+2T)(3M-5T) 6. Factor: 25-y^2 Answer: -(y+5)(y-5) 7. Factor: h^3+9h Answer: h(h^2+9) 8. Factor: a^3+64 (a+4)(a^2-4a+16) 9.Factor: 9y^3-4y Answer: y(3y+2)(3y-2) 10. Factor: S^2+9 Answer: (S+3)(S-3)

OpenStudy (dumbcow):

only #3 and #10 are wrong

OpenStudy (savannah_noelle):

For number 10 would the answer be Prime?

OpenStudy (dumbcow):

yes same for #3

OpenStudy (savannah_noelle):

Okay thanks so much, can you check 10 more? I have one more factoring work sheet

OpenStudy (dumbcow):

ok you know you can check these yourself by just multiplying it out and comparing to original expression

OpenStudy (savannah_noelle):

11. Factor: m^3-8m^2+m-8 Answer: (m^2+1)(m-8) 12. Factor y^3+y-6y^2-6 Answer: (y-6)(y^2+1) 13. Solve: 5x^2+6x+8=0 Answer: No solution 14. Solve: 3x^2-19x+28=0 Answer: ? 15. Solve: (4x-1)(x+9)=0 Answer: x=3, x=-5 16. Solve: (x+3)(x-3)=0 Answer: x=-3, x=3 17. Solve: (4x-1)(x+9)=0 Answer: X=1/4 X==9 18. Solve: 2x^2+5x=3 Answer: x=0 x=-1

OpenStudy (usukidoll):

15. Solve: (4x-1)(x+9)=0 Answer: x=3, x=-5 wrongggggggggggg. solve x-9=0 and then solve for 4x-1=0

OpenStudy (dumbcow):

looks good except for 14,15,18

OpenStudy (savannah_noelle):

@dumbcow Can you show me how to do them?

OpenStudy (savannah_noelle):

Also for number 15 I meant to put the question as : X^2-2x-15=0 Answer: x=3, x=-5

OpenStudy (usukidoll):

oooooh that makes better sense. but the signs for the answers are switched (x+3)(x-5) = 0

OpenStudy (dumbcow):

#14 (3x - __ )(x - __) you need 2 numbers that multiply to 28 ---> (2,14) or (4,7) try a pair and check to see if "x" term is -19 (3x - 2) (x-14) ----> -2x -14(3x) = -44x try again (3x - 7)(x-4) ----> -7x -4(3x) = -19x

OpenStudy (savannah_noelle):

@UsukiDoll Thanks, yeah I had accidentally types a different question. @dumbcow Oh okay thank you! :)

OpenStudy (dumbcow):

18) 2x^2 +5x -3 = 0 (2x + ___) (x - __) 2 numbers must multiply to -3 --> (1,-3) or (-1,3) (2x +1)(x -3) check x term ---> x -3(2x) = -5x we need positive 5 , switch the signs (2x -1)(x +3)

OpenStudy (savannah_noelle):

Thanks again! :)

OpenStudy (dumbcow):

yw